OpenStudy (anonymous):

Use the cubic model y=5a^3-2a^+a-45 to find the value of y when x=4

OpenStudy (anonymous):

y = 5(4)^3 - 2(4)^2 + 4 - 45 = 320 - 32 - 41 Do that expression and you should get your answer

OpenStudy (anonymous):

is it 247?

OpenStudy (anonymous):

Simply substitute (plug in) 4 for x in the expression and evaluate it to find y

OpenStudy (anonymous):

@DribleDaVaca19 Correct!

OpenStudy (anonymous):

oh okay i get it now thank you!

OpenStudy (anonymous):

No problem!